OpenStudy (anonymous):

Which pair of equations does NOT have a common solution? A) x + y = -14x - 3y = 24 B) 4x + 6y = 12 6x + 9y = 12 C) 2x - 3y = -4 2x + y = 4 D) 5x - 4y = 9 x - 2y = -3

OpenStudy (anonymous):

can somebody please help me..... ;P I'm having lots ah trouble with me.........

OpenStudy (anonymous):

B is the correct answer... BUT I GUESSSED. N i WOULD REALLY like to know why?? @TuringTest

OpenStudy (anonymous):

in b , multiply firs eq. by 3/2 u shoul get right hand side of the seconfd eq. as 18 but it is 12 so no common soln.

OpenStudy (anonymous):

Didn't I show you how to do this before? 4x + 6y = 12 6x + 9y = 12 -3(4x + 6y = 12) 2(6x + 9y = 12) -12x + -18y = -36 12x + 18y = 24 0 = -12 See? The variable disappeared and the numbers are different.

OpenStudy (anonymous):

 That's how the two lines look like. They're parallel.